OpenStudy (anonymous):

Solve the inequality. |r + 3| ≥ 7

OpenStudy (campbell_st):

2 solutions \[r + 3 \ge 7 ==========> r \ge 4\] and \[-(r +3) \ge 7 =======> -r \pm 10 ==========> r \le 10\]

OpenStudy (anonymous):

on a side note, those solution sets should be joined by OR.

OpenStudy (anonymous):

check this ,LoL its seem both of u are in the same school xD http://openstudy.com/study#/updates/4fc91170e4b0c6963ad36e7a

OpenStudy (anonymous):

Oh my ghosh, yess we are most likely !